# Optimal quantum simulation of linear non-unitary dynamics

### v1

@@ 5.1 Multiplexed time-evolution operator
Lemma 17 constructs the multiplexed block-encoding $\textsc{Mul}=\sum_{j=-R/h}^{R/h}\lvert j\rangle\langle j\rvert\otimes\textsc{Be}\left[\frac{hjL+H}{\alpha}\right]$, where $\alpha=R\alpha_{L}+\alpha_{H}$ and $R/h\in\mathbb{Z}_{+}$ using one controlled query to the block-encodings of $L$ and $H$, and $\mathcal{O}(\log\frac{R}{h})$ two-qubit gates. By querying controlled-Mul $\mathcal{O}(\alpha t+\log\frac{1}{\epsilon})$ times, then for each $j$, Hamiltonian simulation Lemma 16 block-encodes $\textsc{Be}[U_{j}]\in\mathbb{C}^{2^{m+1}\times 2^{m+1}}$ an operator $U_{j}$ such that $\|U_{j}-e^{-i(hjL+H)t}\|\leq\epsilon$. Hence, we choose
@@ 5.2 Superposition of kernel function amplitudes
    which is block-encoded using $\mathcal{O}(\log\frac{R}{h})$ two-qubit gates, where $U_{\text{r}}$ is from Equation 115, $\textsc{Swap}_{c,d}$ swaps registers $\lvert\cdot\rangle_{c}\lvert\cdot\rangle_{d}$ and $\textsc{Z}_{\text{abs}}$ is the Pauli $Z$ gate on register $\lvert\cdot\rangle_{\text{abs}}$.
@@
    Lemma 20 defines a polynomial $p_{n}$ of degree $n=\mathcal{O}(R\log\frac{R}{\epsilon_{\text{poly}}})$ that approximates $g_{2}$ to error $\epsilon_{k/R}=p_{n}(k/R)-g_{2}(k;\gamma)$ where the error satisfies $\forall x\in[-1,1],\;|\epsilon_{x}|\leq\epsilon_{\text{poly}}$. As $|g_{2}(k;\gamma)|\leq 1$, we may always rescale the polynomial by a small factor $\frac{1}{1+\epsilon_{\text{poly}}}$ without changing the asymptotic complexity $n$ such that $\forall_{x\in[-1,1]}|p_{n}(x)|\leq 1$. Using quantum singular value transformations [GSLW19], we may block-encode $\textsc{Be}[p_{n}[D]]\in\mathbb{C}^{2^{m}\times 2^{m}}$, where $m=2\log\frac{R}{h}+\mathcal{O}(1)$.
@@
We now bound the error of $\lvert\psi^{\prime}_{R,h,\gamma}\rangle$. Let $a_{j}=g_{2}(\frac{hj}{R})$ and $b_{j}=p_{n}(hj/R)$. Then by Lemma 19,
@@
Using the sub-multiplicative property of norms, $\max_{z\in\mathcal{E}_{\rho}}|g_{2}(z;\gamma)|\leq\frac{1}{\sqrt{1-y^{2}}}$. Now rescaling with $z\rightarrow Rz$, $\max_{z\in\mathcal{E}_{\rho}}|g_{2}(Rz;\gamma)|\leq\frac{1}{\sqrt{1-(Ry)^{2}}}=M$, for $y\in(0,1/R)$. Let us choose $y=\frac{1}{2R}$, then $M=\sqrt{4/3}$ and $\rho=1+\frac{1}{2R}+\mathcal{O}(R^{-2})$. Then Lemma 21 on Chebyshev truncation states that the degree-$n$ approximation error is at most
@@ 6 Constant factor optimization
The query complexity of the LCHS is characterized by the normalization constant and the truncation radius. In this section, we numerically optimize the query complexity cost prefactor $\mathrm{cost}=\alpha_{\hat{f}_{j,y},R}\cdot R$ with respect to $R,\gamma,c,j$ for a more general family of kernel functions $\hat{f}_{j,y}$ Equation 6. As the choice of numerical quadrature does not affect query complexity, it suffices to to the error of LCHS directly by Theorem 1. For any target $\epsilon$, the best possible query complexity prefactor for this family of kernel functions may be obtained by numerically minimizing the product
@@
For practical implementation, one should use the numerically optimized LCHS parameters. In Figure 1, we plot this prefactor resulting from solving Equation 162 with $j=2$ and $y=1$ fixed by gradient descent from the parameters proposed in Theorem 2 and find that the analytically derived upper bound is loose by only a factor <1.3<1.3 in the small $\epsilon$ limit. Note that the difference between $\alpha_{\hat{f},R}$ and $\alpha_{\hat{f}}$ in Figure 1 is visually indistinguishable. We also tabulate Table 3 and plot the fits Figure 4 of minimum cost solutions that additionally optimize over $j\geq 1$ and $y>0$. Compared to the $\hat{f}_{2,1}$ case, we find roughly a $25\%$ improvement in the prefactor for $\epsilon\in[10^{-10},10^{-1}]$.
@@
Table 3: For each target error $\epsilon$, we minimize the LCHS cost prefactor $\alpha_{\hat{f}_{j,y},R}\cdot R$ using the generalized kernel functions $\hat{f}_{j,y}$ Equation 6 by solving the optimization program Equation 162 by gradient descent. The first ten row correspond to optimized solutions for $\hat{f}_{2}=\hat{f}_{2,1}$, and the next ten rows allow optimization of $\hat{f}_{j,y}$ over $j$ and $y$ and we find that the optimized solutions have arbitrarily large $\gamma$ which we round to $\infty$. The parameter $y_{0}$ does not affect the definition of the kernel function but is needed for the interested reader to verify the calculations in this table.
Figure 4: Parameter choices of numerically optimized LCHS from Table 3 together with fits to points at $\epsilon\leq 10^{-6}$ for the kernel functions (left) $\hat{f}_{2}(k;\gamma,c)$ and (right) $\hat{f}_{j,y}(k;\gamma,c)$.
@@ 7 Query lower bounds
In this section, we present a $\Omega(u_{\text{lchs}}\alpha_{A}t)$ lower bound Theorem 5 on the query complexity for solving differential equations, and lower bounds on the query complexity of block-encoding $\textsc{Be}[e^{-At}/\alpha_{\hat{f}}$] that suggest our method is optimal in all parameters and cannot be improved by more than a constant factor of $3$. As LCHS block-encodes $e^{-At}$ for non-Hermitian $A=L+iH$, where $L$ and $H$ are Hermitian, any lower bound on block-encoding the special case $e^{-Lt}$ or $e^{-iHt}$ is a lower bound on LCHS. The case $e^{-iHt}$ will not give useful lower bounds on LCHS as it corresponds to the Hamiltonian simulation problem, and the query complexity to the block-encoding $\textsc{Be}[H/\alpha_{H}]$ for $t=\Omega(\log\frac{1}{\epsilon})$ is the optimal value $\alpha_{H}t+\Theta(\log\frac{1}{\epsilon})$ [LC17b, LC19, BMPW24]. Hence, we focus on the case $e^{-Lt}$, which implements imaginary time evolution [SGS+25] and Gibbs sampling [CS17]. Although a $\Omega(\alpha_{L}t)$ lower bound on imaginary time evolution can be proven in a variety of ways [SGS+25], there is no known lower bound that is multiplicative in $\alpha_{L}t$ and $\log\frac{1}{\epsilon}$, which is needed to prove the optimality of our LCHS.
@@
Last we note that, since $N$ and $K$ are independent, so are the norm of the evolved state and evolution time. For example, we can set $K$ and $N$ such that the norm $\sim 1/(2\delta\sqrt{N})=1/(2N^{1/K})$ remains constant, while $t_{N,K}$ still grows with $\sqrt{N}$. This justifies the use of two independent integers, so that the lower bound has an explicit dependence in the norm of the state and evolution time. Additionally, multiplicative scaling in $\alpha_{A}t$ with a tunable factor of any integer $\alpha_{A}=d>0$ can be included by taking the tensor product of $L_{N,K}$ with the $d\times d$ matrix where all entries are unity, which reproduces the exact same evolution in $t/d$ time with the uniform superposition ancilla state [BCC+14].
@@
For general classes of functions, such as $C^{l}$-smooth functions, or analytic functions, the optimal scaling of $n$ is known [Tre19]. However, the optimal scaling when $f$ is defined only piecewise on sub-intervals with a uniform boundedness constraint like $\max_{x\in[-1,1]}|p_{n}(x)|\leq\alpha$ in Equation 182 is poorly understood. We know of only two optimal results for the case $\alpha=\Theta(1)$ – the sign function [EY11] and the inverse function [GSLW19], both of which are characterized by the maximum gradient $\tau$ between two sub-intervals and are approximated by polynomials of optimal degree $\Theta(\tau\log\frac{1}{\epsilon})$. Similarly, we expect the optimal degree for Equation 182 to be $\Theta(\tau\log\frac{1}{\epsilon})$.
Using well-known numerical techniques [HK93], we may rewrite Equation 182 as a semi-infinite program that we discretize into some $K$ points. Any real polynomial of degree $n$ is uniquely characterized by the coefficients $\vec{a}\in\mathbb{R}^{n+1}$ in the Chebyshev expansion
@@
Let $\vec{p}_{n,\text{right}}$ be the subsets of $\vec{p}_{n}$ evaluated at all points $x\geq 0$. Then Equation 182 becomes the $K\rightarrow\infty$ limit of a linear program over $\vec{a}$, which for finite $K$, is solvable in polynomial time.
@@
For any finite $K$, the constraints in Equation 186 are a subset of those in Equation 182. Hence the degree $n$ that achieves a target error $\epsilon$ will be a lower bound on the degree computed by the semi-infinite program. In practice, one should choose $K\gg n$, which will still lead to solutions $p_{n}$ where some points exceed the desired error tolerance, but only by a small multiplicative constant close to unity. As the polynomial of best approximation is characterized by the equioscillation theorem [Tre19], one may iteratively refine the solution of Equation 186 by finding local maxima in error that do not satisfy Equation 186 and adding those points to $\mathcal{X}$. This combined approach augments the Remez exchange algorithm to automatically handle cases of multiple sub-intervals or discontinuous non-uniform error targets where the sign of the error $p_{n}-f$ may not be strictly alternating.

### v2

@@ 5.1 Multiplexed time-evolution operator
Lemma˜18 constructs the multiplexed block-encoding $\textsc{Mul}=\sum_{j=-R/h}^{R/h}\lvert j\rangle\langle j\rvert\otimes\textsc{Be}\left[\frac{hjL+H}{\alpha}\right]$, where $\alpha=R\alpha_{L}+\alpha_{H}$ and $R/h\in\mathbb{Z}_{+}$ using one controlled query to the block-encodings of $L$ and $H$, and $\mathcal{O}(\log\frac{R}{h})$ two-qubit gates. By querying controlled-Mul $\mathcal{O}(\alpha t+\log\frac{1}{\epsilon})$ times, then for each $j$, Hamiltonian simulation Lemma˜17 block-encodes $\textsc{Be}[U_{j}]\in\mathbb{C}^{2^{m+1}\times 2^{m+1}}$ an operator $U_{j}$ such that $\|U_{j}-e^{-i(hjL+H)t}\|\leq\epsilon$. Hence, we choose
@@ 5.2 Superposition of kernel function amplitudes
    which is block-encoded using $\mathcal{O}(\log\frac{R}{h})$ two-qubit gates, where $U_{\text{r}}$ is from Equation˜118, $\textsc{Swap}_{c,d}$ swaps registers $\lvert\cdot\rangle_{c}\lvert\cdot\rangle_{d}$ and $\textsc{Z}_{\text{abs}}$ is the Pauli $Z$ gate on register $\lvert\cdot\rangle_{\text{abs}}$.
@@
    Lemma˜21 defines a polynomial $p_{n}$ of degree $n=\mathcal{O}(R\log\frac{R}{\epsilon_{\text{poly}}})$ that approximates $g_{2}$ to error $\epsilon_{k/R}=p_{n}(k/R)-g_{2}(k;\gamma)$ where the error satisfies $\forall x\in[-1,1],\;|\epsilon_{x}|\leq\epsilon_{\text{poly}}$. As $|g_{2}(k;\gamma)|\leq 1$, we may always rescale the polynomial by a small factor $\frac{1}{1+\epsilon_{\text{poly}}}$ without changing the asymptotic complexity $n$ such that $\forall_{x\in[-1,1]}|p_{n}(x)|\leq 1$. Using quantum singular value transformations [GSLW19], we may block-encode $\textsc{Be}[p_{n}[D]]\in\mathbb{C}^{2^{m}\times 2^{m}}$, where $m=2\log\frac{R}{h}+\mathcal{O}(1)$.
@@
We now bound the error of $\lvert\psi^{\prime}_{R,h,\gamma}\rangle$. Let $a_{j}=g_{2}(\frac{hj}{R})$ and $b_{j}=p_{n}(hj/R)$. Then by Lemma˜20,
@@
Using the sub-multiplicative property of norms, $\max_{z\in\mathcal{E}_{\rho}}|g_{2}(z;\gamma)|\leq\frac{1}{\sqrt{1-y^{2}}}$. Now rescaling with $z\rightarrow Rz$, $\max_{z\in\mathcal{E}_{\rho}}|g_{2}(Rz;\gamma)|\leq\frac{1}{\sqrt{1-(Ry)^{2}}}=M$, for $y\in(0,1/R)$. Let us choose $y=\frac{1}{2R}$, then $M=\sqrt{4/3}$ and $\rho=1+\frac{1}{2R}+\mathcal{O}(R^{-2})$. Then Lemma˜22 on Chebyshev truncation states that the degree-$n$ approximation error is at most
@@ 6 Constant factor optimization
The query complexity of the LCHS is characterized by the normalization constant and the truncation radius. In this section, we numerically optimize the query complexity cost prefactor $\mathrm{cost}=\alpha_{\hat{f}_{j,y},R}\cdot R$ with respect to $R,\gamma,c,j$ for a more general family of kernel functions $\hat{f}_{j,y}$ Equation˜6. As the choice of numerical quadrature does not affect query complexity, it suffices to compute the error of LCHS directly by Theorem˜1. For any target $\epsilon$, the best possible query complexity prefactor for this family of kernel functions may be obtained by numerically minimizing the product
@@
For practical implementation, one should use the numerically optimized LCHS parameters. In Figure˜1, we plot this prefactor resulting from solving Equation˜165 with $j=2$ and $y=1$ fixed by gradient descent from the parameters proposed in Theorem˜2 and find that the analytically derived upper bound is loose by only a factor <1.3<1.3 in the small $\epsilon$ limit. Note that the difference between $\alpha_{\hat{f},R}$ and $\alpha_{\hat{f}}$ in Figure˜1 is visually indistinguishable. We also tabulate Table˜3 and plot the fits Figure˜4 of minimum cost solutions that additionally optimize over $j\geq 1$ and $y>0$. Compared to the $\hat{f}_{2,1}$ case, we find roughly a $25\%$ improvement in the prefactor for $\epsilon\in[10^{-10},10^{-1}]$.
@@
Table 3: For each target error $\epsilon$, we minimize the LCHS cost prefactor $\alpha_{\hat{f}_{j,y},R}\cdot R$ using the generalized kernel functions $\hat{f}_{j,y}$ Equation˜6 by solving the optimization program Equation˜165 by gradient descent. The first ten row correspond to optimized solutions for $\hat{f}_{2}=\hat{f}_{2,1}$, and the next ten rows allow optimization of $\hat{f}_{j,y}$ over $j$ and $y$ and we find that the optimized solutions have arbitrarily large $\gamma$ which we round to $\infty$. The parameter $y_{0}$ does not affect the definition of the kernel function but is needed for the interested reader to verify the calculations in this table.
Figure 4: Parameter choices of numerically optimized LCHS from Table˜3 together with fits to points at $\epsilon\leq 10^{-6}$ for the kernel functions (left) $\hat{f}_{2}(k;\gamma,c)$ and (right) $\hat{f}_{j,y}(k;\gamma,c)$.
@@ 7 Query lower bounds
In this section, we present a $\Omega(u_{\text{lchs}}\alpha_{A}t)$ lower bound Theorem˜5 on the query complexity for solving differential equations, and lower bounds on the query complexity of block-encoding $\textsc{Be}[e^{-At}/\alpha_{\hat{f}}$] that suggest our method is optimal in all parameters and cannot be improved by more than a constant factor of $3$. As LCHS block-encodes $e^{-At}$ for non-Hermitian $A=L+iH$, where $L$ and $H$ are Hermitian, any lower bound on block-encoding the special case $e^{-Lt}$ or $e^{-iHt}$ is a lower bound on LCHS. The case $e^{-iHt}$ will not give useful lower bounds on LCHS as it corresponds to the Hamiltonian simulation problem, and the query complexity to the block-encoding $\textsc{Be}[H/\alpha_{H}]$ for $t=\Omega(\log\frac{1}{\epsilon})$ is the optimal value $\alpha_{H}t+\Theta(\log\frac{1}{\epsilon})$ [LC17b, LC19, BMPW24]. Hence, we focus on the case $e^{-Lt}$, which implements imaginary time evolution [SGS+25] and Gibbs sampling [CS17]. Although a $\Omega(\alpha_{L}t)$ lower bound on imaginary time evolution can be proven in a variety of ways [SGS+25], there is no known lower bound that is multiplicative in $\alpha_{L}t$ and $\log\frac{1}{\epsilon}$, which is needed to prove the optimality of our LCHS.
@@
Lastly, we note that, since $N$ and $K$ are independent, so are the norm of the evolved state and evolution time. For example, we can set $K$ and $N$ such that the norm $\sim 1/(2\delta\sqrt{N})=1/(2N^{1/K})$ remains constant, while $t_{N,K}$ still grows with $\sqrt{N}$. This justifies the use of two independent integers, so that the lower bound has an explicit dependence in the norm of the state and evolution time. Additionally, multiplicative scaling in $\alpha_{A}t$ with a tunable factor of any integer $\alpha_{A}=d>0$ can be included by taking the tensor product of $L_{N,K}$ with the $d\times d$ matrix where all entries are unity, which reproduces the exact same evolution in $t/d$ time with the uniform superposition ancilla state [BCC+14].
@@
For general classes of functions, such as $C^{l}$-smooth functions, or analytic functions, the optimal scaling of $n$ is known [Tre19]. However, the optimal scaling when $f$ is defined only piecewise on sub-intervals with a uniform boundedness constraint like $\max_{x\in[-1,1]}|p_{n}(x)|\leq\alpha$ in Equation˜185 is poorly understood. We know of only two optimal results for the case $\alpha=\Theta(1)$ – the sign function [EY11] and the inverse function [GSLW19], both of which are characterized by the maximum gradient $\tau$ between two sub-intervals and are approximated by polynomials of optimal degree $\Theta(\tau\log\frac{1}{\epsilon})$. Similarly, we expect the optimal degree for Equation˜185 to be $\Theta(\tau\log\frac{1}{\epsilon})$.
Using well-known numerical techniques [HK93], we may rewrite Equation˜185 as a semi-infinite program that we discretize into some $K$ points. Any real polynomial of degree $n$ is uniquely characterized by the coefficients $\vec{a}\in\mathbb{R}^{n+1}$ in the Chebyshev expansion
@@
Let $\vec{p}_{n,\text{right}}$ be the subsets of $\vec{p}_{n}$ evaluated at all points $x\geq 0$. Then Equation˜185 becomes the $K\rightarrow\infty$ limit of a linear program over $\vec{a}$, which, for finite $K$, is solvable in polynomial time.
@@
For any finite $K$, the constraints in Equation˜189 are a subset of those in Equation˜185. Hence the degree $n$ that achieves a target error $\epsilon$ will be a lower bound on the degree computed by the semi-infinite program. In practice, one should choose $K\gg n$, which will still lead to solutions $p_{n}$ where some points exceed the desired error tolerance, but only by a small multiplicative constant close to unity. As the polynomial of best approximation is characterized by the equioscillation theorem [Tre19], one may iteratively refine the solution of Equation˜189 by finding local maxima in error that do not satisfy Equation˜189 and adding those points to $\mathcal{X}$. This combined approach augments the Remez exchange algorithm to automatically handle cases of multiple sub-intervals or discontinuous non-uniform error targets where the sign of the error $p_{n}-f$ may not be strictly alternating.